what is equation of line which is equally inclined to the co ordinate axes and passing through pt (-1,-2,4)?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Hemant
 
Equally inclined means
 
Direction ratio = (1,1,1)
 
Hence
equation of line=
x+1/1 = y +2/1 = y -4/1
 
 
 
Regards
Arun (askIITians forum expert)